# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ 3.2 The cooling and fragmentation of sub-clouds
Figure 5: The cooling diagram of sub-clouds. Black contours show the loci of equaling cooling and free-fall time scales, $t_{\rm cool,sc}=t_{\rm ff,sc}$, at different metallicity (see §3.2 for details). The vertical shaded region indicates the regime where atomic cooling is effective. The horizontal shaded region above the ‘feedback-free’ threshold, $n_{\rm snf}$, indicates the regime where star formation in sub-clouds is free of (internal) supernova feedback. A sub-cloud reaching a density $n_{\rm sc}\geqslant n_{\rm snf}$ thus hosts a GC (see §3.5 for details).
@@
The initial temperature, $10^{4}\,{\rm K}$, set by rapid atomic cooling at the formation of sub-clouds, is much higher than that required by star formation. Thus, other cooling channels, such as those given by metal, dust and molecules, determine the subsequent evolution of sub-clouds until the trigger of star formation.
@@
[…]
@@
where $Z=0.02\,Z_{\odot}\,Z_{0.02}$; $n_{\rm sc}=10^{3.5}\,{\rm cm}^{-3}n_{\rm sc,3.5}$; $T=10^{4}\,{\rm K}\,T_{4}$; a molecular weight $\mu=1.2$ is used; $C=\langle n^{2}\rangle/\langle n\rangle^{2}$ is the clumpiness factor within the sub-cloud. The free-fall time scale depends only on the matter density:
@@
The global collapse of a sub-cloud in the metal-poor condition has been demonstrated by the simulations of Fernandez & Bryan (2018) for individual sub-clouds. They found that, with a moderate UV heating background, a sub-cloud with $M_{\rm sc}=10^{6}{\rm M_{\odot}}$ and $Z=0.01Z_{\odot}$ can collapse globally to form a dense, fragmented core with a size of 5absent5$\approx$ 5≈ 510101010 pc, while the low-density, smooth envelop does not fragment (see their figure 6). A detailed analysis based on the cooling diagram shows that this delayed fragmentation is due to inefficient cooling of metal-poor gas (see their figures 2 and 7). A similar conclusion can be found in the simulations of Grudić et al. (2021) for individual sub-clouds, who found that the star formation efficiency on sub-cloud scales increases with increasing sub-cloud surface density, and that the fraction of stars formed in bound clusters is higher in their low-metallicity runs.
Fig. 6 shows the distributions of sub-cloud density at formation (shaded areas, denoted as $n_{\rm sc}$) and at fragmentation (solid lines, denoted as $n_{\rm sc}^{\prime}$). Details of the modeling steps can be found in §4.1, while here we only focus on the general trend in the sub-cloud density. At $10^{3.5}\,{\rm cm}^{-3}$, the lower limit for a sub-cloud to qualify as (the progenitor of) a globular cluster, only $<10\%$ of sub-clouds are massive enough to form globular clusters, mainly due to the bottom-heavy nature of the sub-cloud mass function (Eq. 30). A significant fraction of metal-poor sub-clouds with low initial density can contract to higher density to reach the condition for globular cluster formation, as clearly suggested by comparing the blue histograms for $n_{\rm sc}$ and $n_{\rm sc}^{\prime}$. The effects of sub-cloud contraction appear to be more significant for low-mass halos, as their metallicity is lower. In contrast, metal-rich sub-clouds hosting globular clusters can fragment immediately after formation. The difference in cooling and fragmentation between the two populations of sub-clouds turns out to be the key to understanding the two channels of GC formation and the observed bimodality in the GC population, as to be discussed in §3.5.
@@ 3.3 Destruction of sub-clouds by external feedback
Once formed, sub-clouds are subject to various external feedback effects produced by other stars in the galaxy. One of the main feedback sources is the shock wave associated with stellar or SN winds. Depending on the state of the gaseous structure that is affected by the wind, the effect of such feedback can be either negative or positive. For dense sub-clouds where cooling is effective and an isothermal state can be maintained, shocks are to compress the gas. On the other hand, for the diffuse component or low-density sub-clouds where gas cannot cool effectively, shocks may disperse gas and disrupt sub-clouds (see e.g. §8.3 of Mo et al., 2010; Hopkins et al., 2012; Rosen, 2022). The effect of the SN feedback is also scale-dependent. At the galactic scale, its effect is to reduce the amount of star-forming gas and prevent star formation, as modeled by the $F_{\rm sn}$ factor in §2.3. Consequently, the total number of sub-clouds survived is also reduced. In terms of individual sub-clouds, the high-density ones are compressed to higher density by shocks, and such positive feedback is already described in §3.1. Here we focus on the negative side of the feedback.
As suggested by Dekel et al. (2023), shock waves associated with stellar or SN winds can ablate or destroy sub-clouds. Following their arguments, we use $t_{\rm crush,sc}$, the time scale for the shock to crush a sub-cloud, to estimate the time needed for the shock to destroy the sub-cloud. This time scale is given by Klein et al. (1994):
@@
where $\Phi$ is the star formation rate of the galaxy, and $\eta_{\rm w}$ is the mass loss fraction of stars. Using the Starburst99 library (Leitherer et al., 1999), and assuming a wind-lasting time of $\approx 10\,{\rm Myr}$ and a specific loss rate of $10^{-8}{\rm yr}^{-1}$, Dekel et al. (2023) estimated that $\eta_{\rm w}\approx 0.1$. The wind velocity is estimated to be $10^{3.5}\,{\rm{km\,s^{-1}}}$, which is the typical value of an SN-driven wind before it is coupled with the ISM. The radius of the SGC, as discussed in §2.2, is $R_{\rm sgc}=f_{\rm gas}R_{\rm v}$. For the sub-cloud to survive the wind disruption and proceed to the regime of collapse and fragmentation, the cooling time scale in the turbulence mixing layer around the sub-cloud has to be shorter than the crushing time scale: $t_{\rm cool,sc}<t_{\rm crush,sc}$. Substituting Eq. (42) and (46) into this condition, we obtain the density threshold for a sub-cloud to survive the wind:
@@
where $\eta_{\rm w}=0.1\eta_{\rm w,0.1}$, $R_{\rm v}=25\,{\rm{kpc}}\,R_{\rm v,25}$ and $\Phi=1{\rm M_{\odot}}/{\rm yr}\,\Phi_{1}$. Using the equation of $\dot{M}_{\rm v}$ in Eq. (6), assuming a star formation efficiency $\epsilon_{\rm v}=M_{*}/(f_{\rm B}M_{\rm v})=0.1$, the mass return fraction $R=0.4$ obtained from the Chabrier IMF, the mean $Z(M_{\rm v},z)$ relation in Eq. (28), and $R_{\rm v}$ in Eq. (4), the shielding density can be approximated as
@@
For both the wind and UV radiation, the $\Phi/R_{\rm sgc}^{2}$-dependence in the shielding density suggests that the survival of sub-clouds is preferred in low-mass, gas-rich halos, because
@@
For SNe inside a sub-cloud, the time scale for them to disperse the sub-cloud is too short for cooling to take place. This sets a time window for star formation and a key criterion for the formation of GCs, as we will describe in §3.4 and §3.5, respectively. The combination of these positive and negative effects will become clearer when we join the pieces to construct a complete model in §4.1. The operation of these processes is also demonstrated by high-resolution zoom-in simulations. For example, as shown in figure 3 of Ma et al. (2020) using the FIRE-2 zoom-in simulations, as soon as SNe are triggered by the formation of a massive star cluster, most of the surrounding gas is dispersed, leaving a number of high-density clumps produced by the compression of stellar/SN winds.
@@ 3.4 Star formation in sub-clouds
$E_{\rm w}=e_{\rm w}\eta_{\rm w}\tau_{\rm sf}\left(\frac{Z}{Z_{\odot}}\right)^{0.9}M_{\rm cls}\,.$ (53)
@@
$\epsilon_{\rm w}$ $=\frac{GM_{\rm sc}}{2R_{\rm sc}(Z/Z_{\odot})^{0.9}\tau_{\rm sf}\eta_{\rm w}e_{\rm w}}$
@@
where $M_{\rm sc}=10^{6}{\rm M_{\odot}}\,M_{\rm sc,6}$, $R_{\rm sc}=12\,{\rm pc}\,R_{\rm sc,12}$, $\tau_{\rm sf}=1\,{\rm Myr}\,\tau_{\rm sf,1}$, $\eta_{\rm w}=0.1\,\eta_{\rm w,0.1}$, and $\epsilon_{\rm w,max}\equiv 3.81/(\tau_{\rm sf,1}\eta_{\rm w,0.1})\sim 1$ is the maximum star formation efficiency. At $Z/Z_{\odot}\approx 0.02$, the star formation efficiency reaches unity for a sub-cloud with mass of $10^{6}{\rm M_{\odot}}$ and a size of $12\,{\rm pc}$. Assuming that $\tau_{\rm sf}$ and $\eta_{\rm w}$ are independent of $M_{\rm sc}$, $R_{\rm sc}$ and $Z$, the star formation efficiency scales as
@@
where $L_{*}/M_{*}\right>\approx 10^{3}L_{\odot}/M_{\odot}$, corresponding to the peak emission of a blackbody radiation with $T\approx 40,000\,{\rm K}$ for OB stars. The boosting factor $\beta_{\rm boost}$ takes into account the effect of multiple scattering. Following Marinacci et al. (2019), we set $\beta_{\rm boost}=1+\kappa_{\rm IR}\Sigma_{\rm sc}$, where $\Sigma_{\rm sc}=M_{\rm sc}/(\pi R_{\rm sc}^{2})$ is the surface density of the sub-cloud and $\kappa_{\rm IR}=10(Z/Z_{\odot})\,{\rm cm^{2}\,g^{-1}}$ is the infrared opacity. Equaling the total momentum injection rate, $\dot{p}_{\rm r}M_{\rm cls}$, to the gravity, $GM_{\rm sc}^{2}/R_{\rm sc}^{2}$, we obtain the star formation efficiency $\epsilon_{\rm r}$ for a radiation-regulated sub-cloud, as
@@
At $Z/Z_{\odot}\approx 0.02$, $\epsilon_{\rm r}$ is proportional to $\Sigma_{\rm sc,3.5}$ and thus we have
@@
This efficiency reaches unity at $\Sigma_{\rm sc}\approx 10^{3.2}\,{\rm M_{\odot}}{\rm pc}^{-2}$, comparable to the condition for sub-clouds with $\epsilon_{\rm w}\approx 1$. At high metallicity, the momentum injection from multiple scattering can effectively reduce the star formation.
@@
where $\epsilon_{\rm sf,max}\approx 1$ is the maximum star formation efficiency, which is reached at $M_{\rm sc}\approx 10^{6}{\rm M_{\odot}}$ and $Z\approx 0.02Z_{\odot}$. The values of $\gamma_{\rm sf}$ and $\beta_{\rm sf}$ depend on the dominant channel of stellar feedback. For example, $(\beta_{\rm sf},\gamma_{\rm sf})=(2/3,-0.9)$ in the wind-regulated regime, while $(\beta_{\rm sf},\gamma_{\rm sf})=(1/3,0)$ in the radiation-regulated, metal-poor regime. If these two channels operate in multiplication, $(\beta_{\rm sf},\gamma_{\rm sf})=(2/3+1/3,-0.9)$.
@@
where $f_{\rm r,cls}\lesssim 1$ depends on the density profile of the cluster. For an isothermal sub-cloud, $f_{\rm r,cls}=1/2$, which we find to be able to reproduce observed sizes of young massive star clusters and globular clusters (see §4.3 below).
@@
where we have defined $M_{\rm cls}=10^{6}{\rm M_{\odot}}\,M_{\rm cls,6}$. In both cases, star clusters have roughly constant size, as the dependence on stellar mass and metallicity is weak. This is indeed seen in observations, where star clusters spanning orders of magnitude in masses have a narrow size range of $1$-$10\,{\rm pc}$ (e.g. Krumholz et al., 2019; Brown & Gnedin, 2021).

### v2

@@ 3.2 The cooling and fragmentation of sub-clouds
Figure 5: The cooling diagram of sub-clouds. Black contours show the loci of equaling cooling and free-fall timescales, $t_{\rm cool,sc}=t_{\rm ff,sc}$, at different metallicity (see §3.2 for details). The vertical shaded region indicates the regime above the critical temperature $T_{\rm cool}=10^{4}\,{\rm K}$, where cooling is effective and fragmentation of SGC is expected to occur to form sub-clouds. The horizontal shaded region above the ‘feedback-free’ threshold, $n_{\rm snf}=10^{3.5}\,{\rm cm}^{-3}$, indicates the regime where star formation in sub-clouds is free of (internal) supernova feedback. A sub-cloud reaching a density $n_{\rm sc}\geqslant n_{\rm snf}$ thus hosts a GC (see §3.5 for details).
@@
The initial temperature, $T_{\rm cool}=10^{4}\,{\rm K}$, set by rapid cooling at the formation of sub-clouds, is much higher than that required by star formation. Thus, other cooling channels, such as those given by metal, dust and molecules, determine the subsequent evolution of sub-clouds until the trigger of star formation.
@@
Following Dekel et al. (2023), we use the dominant channel of $[\textsc{Cii}]\,158\,\mu{\rm m}$ line cooling to estimate the cooling time scale for gas at $Z/\,{\rm Z}_{\odot}>10^{-3}$. Using results given by Krumholz (2012), we have
@@
where $Z=0.02\,\,{\rm Z}_{\odot}\,Z_{0.02}$; $n_{\rm sc}=10^{3.5}\,{\rm cm}^{-3}n_{\rm sc,3.5}$; $T=10^{4}\,{\rm K}\,T_{4}$; a molecular weight $\mu=1.2$ is used; $C=\langle n^{2}\rangle/\langle n\rangle^{2}$ is the clumpiness factor within the sub-cloud. The free-fall timescale depends only on the matter density:
@@
The global collapse of a sub-cloud in the metal-poor condition has been demonstrated by the simulations of Fernandez & Bryan (2018) for individual sub-clouds. They found that, with a moderate UV heating background, a sub-cloud with $M_{\rm sc}=10^{6}{\rm M_{\odot}}$ and $Z=0.01\,{\rm Z}_{\odot}$ can collapse globally to form a dense, fragmented core with a size of 5absent5$\approx$ 5≈ 510101010 pc, while the low-density, smooth envelop does not fragment (see their figure 6). A detailed analysis based on the cooling diagram shows that this delayed fragmentation is due to inefficient cooling of metal-poor gas (see their figures 2 and 7). A similar conclusion can be found in the simulations of Grudić et al. (2021) for individual sub-clouds, who found that the star formation efficiency on sub-cloud scales increases with increasing sub-cloud surface density, and that the fraction of stars formed in bound clusters is higher in their low-metallicity runs.
Grey shades in Fig. 6 show the distributions of sub-cloud density at formation ($n_{\rm sc}$), for all sub-clouds formed in the main branches of central subhalos at $z=0$ with different halo masses. Depending on internal properties and environments, sub-clouds can have different fates: some will be destroyed by external feedback, while the others will continue to contract, fragment and form stars at $n_{\rm sc}^{\prime}$. Only the most massive and/or densest sub-clouds can shield themselves from feedback and form sufficient amounts of stars to qualify as (the progenitors of) globular clusters. These will be discussed thoroughly in §§3.3, 3.4 and 3.5, and be synthesized into a coherent model in §4.1. Briefly, the aforementioned conditions, together with supersonic turbulence and metal-poor environment, give rise to two distinct formation channels that lead to two populations of GCs: Pop-I (the metal-rich population) and Pop-II (the metal-poor population).
The red and blue shades in Fig. 6 show the distributions of gas density at formation for sub-clouds that will eventually evolve into Pop-I and Pop-II GCs, respectively, while the solid lines show the distributions of the gas density at fragmentation for the two populations. At $10^{3.5}\,{\rm cm}^{-3}$, the lower limit for a sub-cloud to qualify as (the progenitor of) a globular cluster, only $<10\%$ of sub-clouds are massive enough to form globular clusters, mainly due to the bottom-heavy nature of the sub-cloud mass function (Eq. 30). A significant fraction of metal-poor sub-clouds with low initial density can contract to higher density to reach the condition for globular cluster formation, as clearly suggested by comparing the blue histograms for $n_{\rm sc}$ and $n_{\rm sc}^{\prime}$. The effects of sub-cloud contraction appear to be more significant for low-mass halos, as their metallicity is lower. In contrast, metal-rich sub-clouds hosting globular clusters can fragment immediately after formation. The difference in cooling and fragmentation between the two populations of sub-clouds turns out to be the key to understanding the two channels of GC formation and the observed bimodality in the GC population, as to be discussed in §3.5.
@@ 3.3 Destruction of sub-clouds by external feedback
Once formed, sub-clouds are subject to various external feedback effects produced by other stars in the galaxy. One of the main feedback sources is the shock wave associated with the energy injection from stellar evolution and SNe, which we collectively refer to as ‘wind feedback’. Depending on the state of the gaseous structure that is affected by the wind feedback, the effect of such feedback can be either negative or positive. For dense sub-clouds where cooling is effective and an isothermal state can be maintained, shocks are to compress the gas. On the other hand, for the diffuse component or low-density sub-clouds where gas cannot cool effectively, shocks may disperse gas and disrupt sub-clouds (see e.g. §8.3 of Mo et al., 2010; Hopkins et al., 2012; Rosen, 2022). The effect of the SN feedback is also scale-dependent. At the galactic scale, its effect is to reduce the amount of star-forming gas and prevent star formation, as modeled by the $F_{\rm sn}$ factor in §2.3. Consequently, the total number of sub-clouds survived is also reduced. In terms of individual sub-clouds, the high-density ones are compressed to higher density by shocks, and such positive feedback is already described in §3.1. Here we focus on the negative side of the feedback.
As suggested by Dekel et al. (2023), wind feedback can ablate or destroy sub-clouds. Following their arguments, we use $t_{\rm crush,sc}$, the timescale for the shock to crush a sub-cloud, to estimate the time needed for the shock to destroy the sub-cloud. This timescale is given by Klein et al. (1994):
@@
where $\Phi$ is the star formation rate of the galaxy, and $\eta_{\rm w}$ is the mass loss fraction of stars. Using the Starburst99 library (Leitherer et al., 1999), and assuming a wind-lasting time of $\approx 10\,{\rm Myr}$ and a specific loss rate of $10^{-8}{\rm yr}^{-1}$, Dekel et al. (2023) estimated that $\eta_{\rm w}\approx 0.1$. The wind velocity is estimated to be $10^{3.5}\,{\rm{km\,s^{-1}}}$, which is the typical value of an SN-driven wind before it is coupled with the ISM. The radius of the SGC, as discussed in §2.2, is $R_{\rm sgc}=f_{\rm gas}R_{\rm v}$. For the sub-cloud to survive the wind disruption and proceed to the regime of collapse and fragmentation, the cooling timescale in the turbulence mixing layer around the sub-cloud has to be shorter than the crushing timescale: $t_{\rm cool,sc}<t_{\rm crush,sc}$. Substituting Eq. (42) and (46) into this condition, we obtain the density threshold for a sub-cloud to survive the wind feedback:
@@
where $\eta_{\rm w}=0.1\eta_{\rm w,0.1}$, $R_{\rm v}=25\,{\rm{kpc}}\,R_{\rm v,25}$ and $\Phi=1{\rm M_{\odot}}/{\rm yr}\,\Phi_{1}$. Using the equation of $\dot{M}_{\rm v}$ in Eq. (6), assuming a star formation efficiency $\epsilon_{\rm v}=M_{*}/(f_{\rm B}M_{\rm v})=0.1$, a mass returning fraction given by Eq. (17), the mean $Z(M_{\rm v},z)$ relation in Eq. (28), and $R_{\rm v}$ in Eq. (4), the shielding density can be approximated as
@@
For both the wind feedback and radiation feedback, the $\Phi/R_{\rm sgc}^{2}$-dependence in the shielding density suggests that the survival of sub-clouds is preferred in low-mass, gas-rich halos, because
@@
For SNe inside a sub-cloud, the timescale for them to disperse the sub-cloud is too short for cooling to take place. This sets a time window for star formation and a key criterion for the formation of GCs, as we will describe in §3.4 and §3.5, respectively. The combination of these positive and negative effects will become clearer when we join the pieces to construct a complete model in §4.1. The operation of these processes is also demonstrated by high-resolution zoom-in simulations. For example, as shown in figure 3 of Ma et al. (2020) using the FIRE-2 zoom-in simulations, as soon as SNe are triggered by the formation of a massive star cluster, most of the surrounding gas is dispersed, leaving a number of high-density clumps produced by the compression of wind feedback.
@@ 3.4 Star formation in sub-clouds
Figure 7: Star formation efficiency (SFE) of sub-clouds as a function of their density ($n_{\rm sc}$), adopting wind feedback and $\epsilon_{\rm w,max}=1$. Black lines show the cases with different sub-cloud mass ($M_{\rm sc}$) and a fixed metallicity, $Z=0.02\,{\rm Z}_{\odot}$. Red lines show the cases with different metallicity and a fixed sub-cloud mass, $M_{\rm sc}=10^{6}{\rm M_{\odot}}$. See §3.4 for details.
@@
$E_{\rm w}=e_{\rm w}\eta_{\rm w}\tau_{\rm sf}\left(\frac{Z}{\,{\rm Z}_{\odot}}\right)^{0.9}M_{\rm cls}\,.$ (53)
@@
$\epsilon_{\rm w}$ $=\frac{GM_{\rm sc}}{2R_{\rm sc}(Z/\,{\rm Z}_{\odot})^{0.9}\tau_{\rm sf}\eta_{\rm w}e_{\rm w}}$
@@
where $M_{\rm sc}=10^{6}{\rm M_{\odot}}\,M_{\rm sc,6}$, $R_{\rm sc}=12\,{\rm pc}\,R_{\rm sc,12}$, $\tau_{\rm sf}=1\,{\rm Myr}\,\tau_{\rm sf,1}$, $\eta_{\rm w}=0.1\,\eta_{\rm w,0.1}$, and $\epsilon_{\rm w,max}\equiv 3.81/(\tau_{\rm sf,1}\eta_{\rm w,0.1})\sim 1$ is the maximum star formation efficiency. Fig. 7 shows the star formation efficiency of sub-clouds as a function of their density, regulated by the wind feedback. At $Z/\,{\rm Z}_{\odot}\approx 0.02$, the star formation efficiency reaches unity for a sub-cloud with mass of $10^{6}{\rm M_{\odot}}$ and a size of $12\,{\rm pc}$, indicating that a low-metallicity sub-cloud free of internal supernova feedback is also free of internal wind feedback. A reduction of density or mass, or an increase of metallicity, decreases the resistance of sub-cloud to wind feedback and thus lowers the star formation efficiency. Assuming that $\tau_{\rm sf}$ and $\eta_{\rm w}$ are independent of $M_{\rm sc}$, $R_{\rm sc}$ and $Z$, the star formation efficiency scales as
@@
where $L_{*}/M_{*}\right>\approx 10^{3}L_{\odot}/M_{\odot}$, corresponding to the peak emission of a blackbody radiation with $T\approx 40,000\,{\rm K}$ for OB stars. The boosting factor $\beta_{\rm boost}$ takes into account the effect of multiple scattering. Following Marinacci et al. (2019), we set $\beta_{\rm boost}=1+\kappa_{\rm IR}\Sigma_{\rm sc}$, where $\Sigma_{\rm sc}=M_{\rm sc}/(\pi R_{\rm sc}^{2})$ is the surface density of the sub-cloud and $\kappa_{\rm IR}=10(Z/\,{\rm Z}_{\odot})\,{\rm cm^{2}\,g^{-1}}$ is the infrared opacity. Equaling the total momentum injection rate, $\dot{p}_{\rm r}M_{\rm cls}$, to the gravity, $GM_{\rm sc}^{2}/R_{\rm sc}^{2}$, we obtain the star formation efficiency $\epsilon_{\rm r}$ for a radiation-regulated sub-cloud, as
@@
At $Z/\,{\rm Z}_{\odot}\approx 0.02$, $\epsilon_{\rm r}$ is proportional to $\Sigma_{\rm sc,3.5}$ and thus we have
@@
This efficiency reaches unity at $\Sigma_{\rm sc}\approx 10^{3.2}\,{\rm M_{\odot}}{\rm pc}^{-2}$, comparable to the condition for sub-clouds with $\epsilon_{\rm w}\approx 1$. At high metallicity, the momentum injection from multiple scattering can effectively reduce the star formation. A more general form of Eq. (55) was originally proposed by Fall et al. (2010), and then confirmed in subsequent hydrodynamical simulations and observations. In particular, Grudić et al. (2018) showed that the formula is accurate for a mixture of feedback processes, both momentum-driven and energy-driven, including radiation pressure, photo-ionization, stellar winds, etc. The formula is also consistent with various observations of proto-clusters/sub-clouds in the MW and the Large Magellanic Cloud (LMC; Mok et al., 2021).
@@
where $\epsilon_{\rm sf,max}\approx 1$ is the maximum star formation efficiency, which is reached at $M_{\rm sc}\approx 10^{6}{\rm M_{\odot}}$ and $Z\approx 0.02\,{\rm Z}_{\odot}$. The values of $\gamma_{\rm sf}$ and $\beta_{\rm sf}$ depend on the dominant channel of stellar feedback. For example, $(\beta_{\rm sf},\gamma_{\rm sf})=(2/3,-0.9)$ in the wind-regulated regime, while $(\beta_{\rm sf},\gamma_{\rm sf})=(1/3,0)$ in the radiation-regulated, metal-poor regime. If these two channels operate in multiplication, $(\beta_{\rm sf},\gamma_{\rm sf})=(2/3+1/3,-0.9)$.
@@
where $f_{\rm r,cls}\lesssim 1$ depends on the density profile of the cluster. For an isothermal sub-cloud, $f_{\rm r,cls}=1/2$, which we find to be able to reproduce observed sizes of young massive star clusters and globular clusters (see §4.3 below). Such a similarity between cluster size and sub-cloud size is expected if star formation in sub-cloud follows a scale-free fragmentation process in which the free-fall timescale is the same throughout the sub-cloud.
@@
[…]